IF Freddy received an 83, 76, 70, 86,and 68 on his first 5 math test; what grade does he need to get pn the sixth test to make his average exactly 80
Answer by mananth(16949)

IF Freddy received an 83, 76, 70, 86,and 68 on his first 5 math test; what grade does he need to get pn the sixth test to make his average exactly 80
..
Total scores of 5 math tests = 403
assume he has to get x
average = 80
(403+x)/6 =80
multiply by 6
403 +x = 480
-403
x=480-403
x=77
...
m.ananth@hotmail.ca